a line that passes through the points (-2,4) and (2,-1). what are the rise and the run of the line? at first i got the answer: the run 4, the rise 5, but it wasn't right. now i think it's -4 and -5. am i right?
Answer by stanbon(75887)

Your y values are going from 4 to -1; that is a decrease of 5
Your x values are going from -2 to 2; that is an increase of 4
So the "run" is 4 and the "rise" is -5
The general way of doing this is to subtract one y-value from
the other and then to subtract the CORRESPONDING x-values.
Doing this you could have either of the following:
[-1-4]/(2-(-2)]=-5/4
or you could have:
[4-(-1)[/[-2-2]= 5/(-4)
In both cases the final answer is -5/4
Cheers,
Stan H.